Session: Nov 1997 (session 4)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Maccaferri A at Fri Nov 14 04:39:59 MET 1997) Minor failures
vc07 to vc14 has no power out

HartRAO (Jonathan Quick at Fri Nov 14 04:42:28 MET 1997) Minor failures
Missed first minute due to lack of checkout time. Both scnas 2 and 3 ran out of tape.

Westerbork (tony foley at Fri Nov 14 09:36:19 MET 1997) Minor failures
Only VC1 to VC4 had internal 0.5MHz filters. VC5 to VC7 had external 0.5MHz filters, but the total power levels were too high.

Noto (C.Trigilio at Fri Nov 14 09:39:47 MET 1997) Success

Torun (K.Borkowski at Fri Nov 14 10:36:36 MET 1997) Success
The tape did not move smoothly which may have affected the quality of recording.

Effelsberg (dgraham at Fri Nov 14 13:36:15 MET 1997) Failure
The operator made a mistake which we cannot reconstruct. All lost.

Jodrell (Alastair Gunn at Sun Nov 16 15:09:16 MET 1997) Success
Track 21 dead. No other known problems.

Calibration (Fredrik T Rantakyro at Tue Dec 9 14:56:24 MET 1997) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the NOV97 directory under the name cah1b.antab Read the comments in the file to get further information.